# SciCoQA: Quality Assurance for Scientific Paper–Code Alignment

### v1

@@ F.3 Code Only Ablation Results
GPT-5 $45.7$ $27.2$ 18.5-18.5 $70.4$ $51.7$ 18.7-18.7 $67.1$ $48.4$ 18.7-18.7
GPT-5 Mini $44.4$ $18.5$ 25.9-25.9 $65.7$ $38.7$ 27.0-27.0 $62.8$ $36.0$ 26.8-26.8
Gemini 2.5 Pro $39.5$ $14.8$ 24.7-24.7 $48.3$ $34.5$ 13.8-13.8 $47.1$ $31.9$ 15.2-15.2
GPT-5 Codex $29.6$ $12.3$ 17.3-17.3 $48.7$ $28.9$ 19.8-19.8 $46.2$ $26.7$ 19.5-19.5
GPT-OSS 20B $42.0$ $16.0$ 25.9-25.9 $46.8$ $33.4$ 13.4-13.4 $46.2$ $31.1$ 15.1-15.1
GPT-OSS 120B $35.8$ $23.5$ 12.3-12.3 $44.7$ $37.2$ 7.5-7.5 $43.5$ $35.4$ 8.2-8.2
Gemini 2.5 Flash $35.8$ $16.0$ 19.8-19.8 $40.9$ $29.6$ 11.3-11.3 $40.3$ $27.8$ 12.4-12.4
GPT-5 Nano $19.8$ $8.6$ 11.1-11.1 $28.9$ $19.4$ 9.4-9.4 $27.7$ $18.0$ 9.7-9.7
Average $36.6$ $17.1$ 19.4-19.4 $49.3$ $34.2$ 15.1-15.1 $47.6$ $31.9$ 15.7-15.7
Table 6: Discrepancy recall depending on the input: The P+C column refers to the performance when providing paper and code, C for only providing the code, and $\Delta$ shows the absolute difference between the two.
Appendix G Example Discrepancies in SciCoQA
@@
Paper Omission Algorithm The paper conditions layers using FiLM by mapping the loss-parameter vector λ through two MLPs to obtain σ and μ and then applying the affine transform $\tilde{f}$ = σ ⊙ f + μ; there is no mention of any output activation on σ or μ, and in the style transfer details it explicitly states that σ and μ are computed via affine maps. In the repository, the FiLMBlock applies a sigmoid activation to both μ and σ before using them, thereby constraining both to the (0,1) range. Thus, the code restricts the FiLM coefficients to be positive and bounded, whereas the paper’s description implies unconstrained affine coefficients produced by linear layers. This is a clear deviation between the implementation and the method described in the paper.
@@
Paper Omission Model The paper states that, following GNNExplainer’s setup, a three-layer GNN is first trained and then explained, but it does not mention any batch normalization in this backbone network. In the public code for node-classification tasks (synthetic datasets), the GCN model enables batch normalization by default (args.bn=True) and applies tf.keras.layers.BatchNormalization after all GCN layers except the last, i.e., after the first and second layers in a 3-layer GCN. This BatchNorm usage is not described in the paper’s model description or experimental setup, making it an undocumented architectural difference between the paper and the released code.
@@
Code Omission Data The paper describes training classifiers using a combination of original MNIST data and CGN-generated counterfactual images. This is explicitly stated in Section 3.2 and reinforced by the MNIST ablation (Appendix A.3) and Table 2 which presents an “Original + CGN” setting. In contrast, the released MNIST training code (mnists/train_classifier.py) only supports training on a single dataset per run; the provided datasets and README usage instruct training on the counterfactual dataset alone (e.g., wildlife_MNIST_counterfactual) without combining with original data. Thus, the code implements CF-only training for MNIST classifiers rather than the paper’s combined original+CF setup, constituting a mismatch in the MNIST training protocol.
@@
Difference Eval The paper specifies “accuracy-based confidence” for a question as the mean accuracy across K sampled responses, i.e., the fraction of correct generations among the K samples. In the code, however, the function calculate_confidence takes the list of per-sample correctness indicators (greedy_scores) and returns max(greedy_scores), which reduces confidence to a binary value: 1 if any sample is correct, 0 otherwise. Moreover, while the code also computes greedy_scores_avg (the intended mean), it does not use that for the confidence fed into the downstream datasets; instead, the binary max value is stored as "confidence" and used for reward/align data. Thus, the implemented confidence deviates from the paper’s definition by using maximization rather than averaging.
Difference Loss The paper defines the second penalty P2 to enforce the admissibility constraints (φ, φ^c) ∈ ADM(c) by penalizing only negative values of ξ(x,y) = c(x,y) − φ(x) − φ^c(y), specifically using min(ξ(x,y), 0)^2. In the code, P2 is implemented in admissable_penalty as torch.mean(torch.clamp(full_xi_vals, min=0)**2), which penalizes positive ξ values (since negative values are clamped to zero). Therefore, the implementation reverses the intended sign: the paper penalizes violations where ξ < 0, while the code penalizes cases where ξ > 0. This directly contradicts the paper’s definition for P2.
Table 7: Examples from the SciCoQA dataset, including discrepancy type and category annotations. All samples are from real data.
Appendix H Example Synthetic Discrepancies in SciCoQA

### v2

@@ C.8 Code Only Ablation Results
GPT-5 $41.3$ $29.3$ 12.0-12.0 $70.0$ $51.0$ 19.0-19.0 $65.8$ $47.9$ 18.0-18.0
GPT-5 Mini $46.7$ $23.9$ 22.8-22.8 $64.3$ $39.6$ 24.7-24.7 $61.7$ $37.3$ 24.4-24.4
Gemini 3.1 Pro $46.7$ $25.0$ 21.7-21.7 $56.4$ $42.0$ 14.4-14.4 $55.0$ $39.5$ 15.4-15.4
GPT-OSS 20B $42.4$ $17.4$ 25.0-25.0 $47.1$ $33.3$ 13.8-13.8 $46.5$ $31.0$ 15.4-15.4
Gemini 2.5 Pro $39.1$ $18.5$ 20.7-20.7 $48.4$ $33.1$ 15.3-15.3 $47.1$ $31.0$ 16.1-16.1
GPT-5 Codex $27.2$ $12.0$ 15.2-15.2 $48.6$ $28.0$ 20.6-20.6 $45.5$ $25.7$ 19.8-19.8
GPT-OSS 120B $41.3$ $22.8$ 18.5-18.5 $44.0$ $34.6$ 9.4-9.4 $43.6$ $32.9$ 10.7-10.7
Gemini 2.5 Flash $34.8$ $17.4$ 17.4-17.4 $41.4$ $28.5$ 12.9-12.9 $40.5$ $26.9$ 13.5-13.5
Average $39.9$ $20.8$ 19.2-19.2 $52.5$ $36.3$ 16.3-16.3 $50.7$ $34.0$ 16.7-16.7
Table 10: Discrepancy recall of the top 8 performing models depending on the input: The P+C column refers to the performance when providing paper and code, C for only providing the code, and $\Delta$ shows the absolute difference between the two.
Appendix D Example Discrepancies in SciCoQA
@@
Paper Omission Algorithm Gemini 3.1
The paper discusses utilizing Feature-wise Linear Modulation (FiLM) to adjust convolutional model features based on the loss weights. For this modulation, it describes using multi-layer perceptrons to compute unconstrained scaling ($\sigma$) and shifting ($\mu$) parameter bounds which are applied as affine transformations to the features. In contrast, the code implements a ‘FiLMBlock‘ that processes the shift (‘mu‘) and scale (‘sigma‘) through a sigmoid activation function prior to applying them to the features. The difference is the undocumented use of an output sigmoid layer on the FiLM parameters in the implementation which restricts the scaling factor to solely shrink the features and the shift translation strictly to positive values compared to the unrestricted parameter bounds discussed in the text.
GPT-5
@@
The paper conditions layers using FiLM by mapping the loss-parameter vector λ through two MLPs to obtain σ and μ and then applying the affine transform $\tilde{f}$ = σ ⊙ f + μ; there is no mention of any output activation on σ or μ, and in the style transfer details it explicitly states that σ and μ are computed via affine maps. In the repository, the FiLMBlock applies a sigmoid activation to both μ and σ before using them, thereby constraining both to the (0,1) range. Thus, the code restricts the FiLM coefficients to be positive and bounded, whereas the paper’s description implies unconstrained affine coefficients produced by linear layers. This is a clear deviation between the implementation and the method described in the paper.
Paper Omission Model Gemini 3.1
The paper states that a three-layer GNN is trained for the node-classification tasks before generating explanations, but omits any description of batch normalization layers being used as part of the architecture. However, the code implementation for node classification models includes batch normalization by default (‘args.bn = True‘). Specifically, in the GCN model code, batch normalization is applied after the first and second GCN layers during the forward pass. This introduces an undocumented architectural choice present in the code but not mentioned in the paper.
GPT-5
@@
The paper states that, following GNNExplainer’s setup, a three-layer GNN is first trained and then explained, but it does not mention any batch normalization in this backbone network. In the public code for node-classification tasks (synthetic datasets), the GCN model enables batch normalization by default (args.bn=True) and applies tf.keras.layers.BatchNormalization after all GCN layers except the last, i.e., after the first and second layers in a 3-layer GCN. This BatchNorm usage is not described in the paper’s model description or experimental setup, making it an undocumented architectural difference between the paper and the released code.
Code Omission Data Gemini 3.1
The paper states that for the MNIST experiments, the "Original + CGN" baseline is trained on a combination of real and counterfactual images. However, the implemented code for training the MNIST classifiers does not combine these datasets. The data loading logic defined in ‘get_tensor_dataloaders‘ exclusively loads the counterfactual tensor dataset when a counterfactual dataset name is evaluated, without appending or mixing in the original training dataset. Therefore, the code implementation trains solely on counterfactual data, lacking the original data indicated by the paper.
GPT-5
@@
The paper describes training classifiers using a combination of original MNIST data and CGN-generated counterfactual images. This is explicitly stated in Section 3.2 and reinforced by the MNIST ablation (Appendix A.3) and Table 2 which presents an “Original + CGN” setting. In contrast, the released MNIST training code (mnists/train_classifier.py) only supports training on a single dataset per run; the provided datasets and README usage instruct training on the counterfactual dataset alone (e.g., wildlife_MNIST_counterfactual) without combining with original data. Thus, the code implements CF-only training for MNIST classifiers rather than the paper’s combined original+CF setup, constituting a mismatch in the MNIST training protocol.
Table 11: Discrepancy ground truth examples from the SciCoQA dataset generated by both Gemini 3.1 Pro and GPT-5, including discrepancy type and category annotations. All samples are from real data.
Type Category Description
Code Omission Training Gemini 3.1
The paper algorithm outlines that within each global aggregation round, each worker performs $\tau$ local update steps on different minibatches and accumulates the computed gradients to represent the overall worker gradient. However, in the code, the ‘federated_worker_process‘ loops over the dataloader, performs a single forward and backward pass, and then immediately unconditionally breaks out of the loop. Additionally, the lines of code previously intended for accumulating gradients (‘accum_grads = add_param_list(accum_grads, worker_grads)‘) are commented out, which means the code effectively enforces $\tau=1$.
GPT-5
@@
Difference Eval Gemini 3.1
The paper describes the accuracy-based confidence score as the mean accuracy of the candidate responses, computed by taking the average of the indicator functions that check if each sampled response matches the ground-truth answer. In the provided code, the confidence score calculation is implemented by returning the maximum value of the sampled correctness scores (‘max(probabilities)‘) instead of their average. This creates a discrepancy where the code leverages a binary indicator reflecting whether at least one correct response was generated (1 if yes, 0 otherwise), rather than the continuous proportion of correct responses specified by the formula in the paper.
GPT-5
The paper specifies “accuracy-based confidence” for a question as the mean accuracy across K sampled responses, i.e., the fraction of correct generations among the K samples. In the code, however, the function calculate_confidence takes the list of per-sample correctness indicators (greedy_scores) and returns max(greedy_scores), which reduces confidence to a binary value: 1 if any sample is correct, 0 otherwise. Moreover, while the code also computes greedy_scores_avg (the intended mean), it does not use that for the confidence fed into the downstream datasets; instead, the binary max value is stored as "confidence" and used for reward/align data. Thus, the implemented confidence deviates from the paper’s definition by using maximization rather than averaging.
Difference Loss Gemini 3.1
The paper defines the $P_{2}$ penalty term as the sum of $\min(\xi(x,y),0)^{2}$, which applies a squared penalty strictly to negative evaluations of $\xi(x,y)$ in order to enforce the provided optimal transport constraints. However, the code implements this term using ‘torch.clamp(full_xi_vals, min=0)**2‘. The PyTorch function with ‘min=0‘ imposes a lower bound of 0, mathematically returning $\max(\xi(x,y),0)^{2}$, which zeroes out the negative values and squares the positive ones. Consequently, the codebase incorrectly penalizes positive values of $\xi(x,y)$ instead of negative values as stated in the methodology.
GPT-5
The paper defines the second penalty P2 to enforce the admissibility constraints (φ, φ^c) ∈ ADM(c) by penalizing only negative values of ξ(x,y) = c(x,y) − φ(x) − φ^c(y), specifically using min(ξ(x,y), 0)^2. In the code, P2 is implemented in admissable_penalty as torch.mean(torch.clamp(full_xi_vals, min=0)**2), which penalizes positive ξ values (since negative values are clamped to zero). Therefore, the implementation reverses the intended sign: the paper penalizes violations where ξ < 0, while the code penalizes cases where ξ > 0. This directly contradicts the paper’s definition for P2.
Table 11: Discrepancy ground truth examples from the SciCoQA dataset generated by both Gemini 3.1 Pro and GPT-5, including discrepancy type and category annotations. All samples are from real data. (Continued)
Appendix E Example Synthetic Discrepancies in SciCoQA
@@
Appendix F Prompts
F.1 GitHub Issue Discrepancy Extraction
PROMPT  | GitHub Issue Discrepancy Extraction You are an assistant that analyzes GitHub issues of scientific codebases. Your primary goal is to determine if the code repository contains any inconsistency, discrepancy, or mismatch between what is described in the paper and implemented in the code. For this, you analyze a GitHub issue and determine whether it reports a concrete discrepancy. A **concrete discrepancy** means the issue clearly describes a mismatch between what is stated in the paper (e.g., formulas, algorithms, hyperparameters, methods, logic, or processes) and what is implemented in the repository’s code. Important: Only label issues as a discrepancy if they point to a _specific, concrete difference_ between paper and code. Do not label general reproducibility problems, missing details, or unrelated bugs as discrepancies. ## Not a discrepancy issue Label as **Not a discrepancy issue** if: - The issue is about anything other than a difference between the paper and the code. - The issue describes reproducibility problems (e.g., different results) but does **not** identify a concrete paper-code mismatch. - The issue is about missing information needed to reproduce results without pointing to a mismatch. - The issue is about bugs or errors unrelated to the paper’s described methods or experiments. ## Discrepancy issue Label as **Discrepancy issue** if: - The issue explicitly reports a mismatch between the paper and the code implementation. - The mismatch can involve hyperparameters, formulas, algorithms, logic, processes, or other settings described in the paper. ## Response Format After analyzing the issue in detail and applying the definitions above, provide your final classification in the structure defined below: ### Final Answer “‘yaml issue_label: <the issue label: "Not a discrepancy issue" or "Discrepancy issue"> “‘ ## Issue {issue}
F.2 GitHub Issue Discrepancy Verification
PROMPT  | GitHub Issue Discrepancy Verification Your task is to verify a paper-code discrepancy described in a GitHub issue. Your goal is to verify whether the discrepancy is valid or not. Follow these steps to verify the discrepancy: 1. Analyze the issue and ensure you understand exactly the claimed discrepancy between the paper and the code. 2. Analyze the paper and the code and understand in detail the relevant paper sections and code files. 3. Using your understanding of paper, code, and the discrepancy in the issue, analyze whether the discrepancy is valid or not. A discrepancy is valid, if the claimed discrepancy actually exists and there is a difference between the paper description and the code. 4. Provide your final judgement whether the reported discrepancy is valid or not, and if so a summary and the relevant paper sections and code files in the format below: “‘yaml is_valid_discrepancy: <yes or no> is_valid_discrepancy_reason: <provide a short explanation for your judgement> discrepancy_summary: <if valid provide the following description, else this should be empty: a summary of the discrepancy between the paper and the code in 3-8 sentences. Your description should contain three parts focusing on the discrepancy: 1) summarize what is described in the paper, 2) summarize what is implemented in the code, and 3) summarize the difference. Do not speculate about the impact.> relevant_paper_sections: - <verbatim quote any parts from the paper that are relevant to the discrepancy> - <if there are multiple relevant parts, paste each of them.> relevant_code_files: - <name any code files that are relevant to the discrepancy by providing the file name> - <if there are multiple relevant code files, paste each of them.> “‘ ## Issue {issue} ## Paper {paper} ## Code {code}
F.3 Reproducibility Paper Discrepancy Extraction
PROMPT  | Reproducibility Paper Discrepancy Extraction You are an assistant that analyzes reproducibility reports of scientific papers. Your primary goal is to detect whether the report identifies any discrepancies between the original paper and the original code repository. ## What counts as a discrepancy - A concrete paper–code discrepancy means the report clearly describes a mismatch between what is stated in the original paper (e.g., formulas, algorithms, logic, methods, processes, or other settings) and what is implemented in the original code repository. - Each distinct mismatch should be reported as a separate item. - Hyperparameter mismatches (e.g., learning rate, batch size, dropout rate) do not count as discrepancies, since these are typically configurable in code repositories. - If the report only describes differences between reproduced results and original results, without identifying a paper–code mismatch, do not list it. - If the report speculates about a possible mismatch (uncertain or ambiguous wording), still list it, but mark it with confindence: low. ## What does not count as a discrepancy - General reproducibility problems (e.g., “we could not match the reported results”). - Missing information in the paper (e.g., “the authors did not specify X”). - Missing implementation in the original code repository (e.g., “the authors did not provide the code for X”). - Bugs or errors in the code that are unrelated to what the paper describes. - Differences between reproduced implementation/results and the original paper. ## Output format Summarize all discrepancies found in the following structure, providing a description of the discrepancy, evidence supporting the discrepancy as verbatim quotes from the reproducibility report, and a confidence estimate of the authors on the reported discrepancy. “‘yaml concrete_paper_code_discrepancies: - description: "<3–8 sentence descriptive summary of the first discrepancy>" evidence: - "<paste any evidence (e.g. a paragraph describing the discrepancy) from the reproducibility report that support the discrepancy.>" - "<If there are multiple evidence, paste each of them.>" confidence: <Estimate of the confidence the authors have on the reported discrepancy. One of: low, medium, high>- … “‘ If no discrepancies are reported, return: “‘yaml concrete_paper_code_discrepancies: [] “‘ Reproducibility Report: {paper}
F.4 Reproducibility Paper Discrepancy Verification
PROMPT  | Reproducibility Paper Discrepancy Verification Your task is to verify a paper-code discrepancy described in a reproducibility report. You will be provided with a summary of the discrepancy according to the report, and potentially multiple quotes from the reproducibility report that support the discrepancy. Your goal is to verify whether the discrepancy is valid or not. Follow these steps to verify the discrepancy: 1. Analyze the description and evidence from the reproducibility report and ensure you understand exactly the claimed discrepancy between the paper and the code. 2. Analyze the paper and the code and understand in detail the relevant paper sections and code files. 3. Using your understanding of paper, code, and the reported discrepancy, analyze whether the discrepancy is valid or not. A discrepancy is valid, if the claimed discrepancy actually exists, i.e. the described difference between paper and code exists. 4. Provide your final judgement whether the reported discrepancy is valid or not, and if so a summary and the relevant paper sections and code files in the format below: “‘yaml is_valid_discrepancy: <yes or no> is_valid_discrepancy_reason: <provide a short explanation for your judgement> discrepancy_summary: <if valid provide the following description, else this should be empty: a summary of the discrepancy between the paper and the code in 3-8 sentences. Your description should contain three parts focusing on the discrepancy: 1) summarize what is described in the paper, 2) summarize what is implemented in the code, and 3) summarize the difference. Do not speculate about the impact.> relevant_paper_sections: - <verbatim quote any parts from the paper that are relevant to the discrepancy> - <if there are multiple relevant parts, paste each of them.> relevant_code_files: - <name any code files that are relevant to the discrepancy by providing the file name> - <if there are multiple relevant code files, paste each of them.> “‘ ## Reproducibility Report with Discrepancy {discrepancy_in_report} ## Paper {paper} ## Code {code}
F.5 Synthetic Discrepancy Generation
Below the prompt for generating synthetic discrepancies for computer science papers. For non-CS paper, we remove the discrepancy type definition and prediction from the prompt.
PROMPT  | Synthetic Discrepancy Generation Your task is to generate 5 realistic paper–code discrepancies by introducing small, conceptually meaningful modifications to the codebase of the provided research paper. Follow these steps to perform the edit: 1. Carefully read and understand both the research paper and the entire code repository provided below. Your goal is to identify the key ideas, methods, and components described in the paper and how they correspond to the implementation in the code. 2. Your changes must adhere to the following constraints: - **Small**: The changes must affect a few lines of code or a short function. It may affect multiple files, but only to the extent necessary to create a coherent and realistic discrepancy. - **Relevance**: The changes must relate directly to a core scientific or algorithmic idea of the paper and would likely have an impact on reproducibility of results and validity of claims. - **Significance**: The changes must introduce a conceptual discrepancy — not a simple hyperparameter, or formatting change, or a change that could be fixed via a simple (command line) argument. - **Scope**: The changes should not all rely on already implemented features of the code base, but also implement new features or introduce modifications. Balance between relying on existing features and implementing new features. - **No Comments**: Do not add comments to the changed code which would easily allow to identify the discrepancy. - **No Bugs**: The introduced discrepancies should not be bugs, i.e., code that could be detected as erroneous by inspecting the code itself. The discrepancy must be related to both the paper and the code to be identified. - **No Implications**: The discrepancies must not rely on anything the paper only implies or assumes, but the paper must clearly conflict with the code after the change, or omit an important concept that is implemented in the changed code. 3. Your changes can be one of the following types. You can create multiple discrepancies of the same or different types. - **Paper Omission**: Modify the code such that it implements a concept or idea that is not described in the paper. - **Code Omission**: Modify the code such that it drops a specific concept or idea that is described in the paper. - **Difference**: Modify the code such that there is a difference between the paper and the code, i.e., the paper describes one thing and the code implements another, e.g., by changing the order of steps, operations, or core logic. 4. Decide which paper-code discrepancies are most appropriate for the given paper, choosing from the following categories. Note, you can create multiple discrepancies of the same or different types. - **Loss**: changes to loss definition or terms - **Algorithm**: changes in the order of steps, operations, or core logic - **Training**: changes to the learning process, schedule, or optimization - **Evaluation**: changes to evaluation logic, metrics, or scripts - **Model**: architectural or initialization changes - **Data**: dataset usage, preprocessing, augmentation, or filtering - **Other**: other types of discrepancies that are not covered by the above categories but appropriate for the given paper. 5. Generate 5 discrepancies in the following strict format: “‘md # Discrepancy 1 - Type: <choose one from: Paper Omission, Code Omission, Difference> - Category: <choose one from: Loss, Algorithm, Training, Evaluation, Model, Data, Other> - Description: <a summary of the discrepancy between the paper and the code in 3-8 sentences. When referring to the code, do not mention ’original’ or ’modified’, but assume the code is published with the modification and discrepancy. Your description should contain three parts focusing on the discrepancy: 1) summarize what is described in the paper, 2) summarize what is implemented in the modified code, and 3) summarize the difference. Do not speculate about the impact.> ## Code Changes <<<<<<<ORIGINAL CODE: <relative/path/to/file.py> (paste the relevant lines of the original code exactly as they appear) ======= (paste the lines of the modified code containing your change that replace the original code) >>>>>>>DISCREPANCY If multiple files are affected, repeat the diff block for each. ## Relevant Paper Sections - <verbatim quote any parts from the paper that are relevant to the discrepancy.> - <if there are multiple relevant parts, paste each of them.> # Discrepancy 2 # Discrepancy 3 “‘ ## Paper {paper} ## Code {code}
F.6 Discrepancy Prediction
PROMPT  | Discrepancy Prediction You are an expert in analyzing scientific papers and their code implementations. Your task is to carefully identify concrete discrepancies between what is described in a paper and what is actually implemented in the code. ## What counts as a discrepancy - A concrete paper–code discrepancy means a mismatch between what is stated in the original paper (e.g., formulas, algorithms, logic, methods, processes, or other settings) and what is implemented in the original code repository. - Each distinct mismatch should be reported as a separate item. ## What does not count as a discrepancy - Missing information in the paper like hyperparameters (e.g., “the authors did not specify X”). - Hyperparameter mismatches (e.g., learning rate, batch size, dropout rate), since these are typically configurable in code repository. - Missing implementation in the original code repository (e.g., “the authors did not provide the code for X”). - Bugs or errors in the code that are unrelated to what the paper describes. ## Output format Provide your findings in the following YAML structure: “‘yaml discrepancies: - <a summary of the discrepancy between the paper and the code in 3-8 sentences. Your description should contain three parts focusing on the discrepancy: 1) summarize what is described in the paper, 2) summarize what is implemented in the code, and 3) summarize the difference. Do not speculate about the impact.> - <if there are multiple discrepancies, put each of them in a separate item.> “‘ ## Paper {paper} ## Code
F.7 Discrepancy Evaluation
PROMPT  | Discrepancy Evaluation Your task is to evaluate whether a reference paper - code discrepancies matches a predicted paper - code discrepancy. Follow these steps: 1. Analyze which part of the paper or code each discrepancy is describing. Extract the core claims and issues from the reference and predicted discrepancies. 2. Analyze whether the core claims are about the same issue, i.e. if they describe the same or different paper-code discrepancies. The two discrepancies might use different wording or one might be more detailed than the other. Focus on whether the issue is the same, even if minor details are different. However, if they describe different issues (even about the same topic or part of the paper or code) they do not match. 3. Provide a brief explanation of your reasoning. ## Reference Paper-Code Discrepancy {reference discrepancy} ## Predicted Paper-Code Discrepancy {predicted discrepancy} ## Answer Format Provide your answer in the following format: “‘yaml core_claim_reference: <core claim from reference discrepancy > core_claim_predicted: <core claim from predicted discrepancy > reasoning: <explanation of why the core claims concern the same issue > match: <yes | no > “‘